OpenStudy (anonymous):

if 0.00531 mol N2O effuses through an orifice in a certain period of time, how many moles of N2O3 would effuse in the same time under the same conditions?

OpenStudy (aaronq):

use grahams law of diffusion

OpenStudy (anonymous):

How though? I know that the rate of two things divided is equal to the inverse square root of the molar masses but how do I apply that?

OpenStudy (aaronq):

when you find the ratio of their rates of diffusion apply that. say the rate of \(\dfrac{N_2O}{N_2O_3}\) is \(\dfrac{2}{1}\), then just make them equal to each other. \(\dfrac{N_2O}{N_2O_3}=\dfrac{2}{1}\rightarrow\dfrac{0.00531mol }{n_{N_2O_3}}=\dfrac{2}{1} \rightarrow n_{N_2O_3}=0.002655 mol\)